OpenStudy (anonymous):

Indicate whether the conjecture is true or false. If false, provide a counterexample.

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

do you think it's true or false?

OpenStudy (anonymous):

I'm not sure how to find out.

jimthompson5910 (jim_thompson5910):

if you square a positive number, then take the square root of that result, what do you get?

OpenStudy (anonymous):

so like 5 squared = 25 so the square root is 5

jimthompson5910 (jim_thompson5910):

how about squaring 7 and then taking the square root of that

OpenStudy (anonymous):

it equals 7

OpenStudy (anonymous):

so it would be true right?

jimthompson5910 (jim_thompson5910):

so for any positive number, squaring it, then taking the square root gives you the same number

jimthompson5910 (jim_thompson5910):

what happens when you start with a negative number?

OpenStudy (anonymous):

it comes out negative

OpenStudy (anonymous):

no wait....i dont think a negative can have a squareroot right?

jimthompson5910 (jim_thompson5910):

say you start with -10 square -10 to get ???

OpenStudy (anonymous):

3.16?

jimthompson5910 (jim_thompson5910):

square, as in multiply it by itself

OpenStudy (anonymous):

opps sorry...it would be -100

jimthompson5910 (jim_thompson5910):

no, -10 squared = (-10) * (-10) = ???

jimthompson5910 (jim_thompson5910):

when we have x^2 and we replace x with -10, then we square all of -10

jimthompson5910 (jim_thompson5910):

x^2 = (-10)^2 = (-10)*(-10) = ???

OpenStudy (anonymous):

im getting confused i swear i squared -10 right...

OpenStudy (anonymous):

nope i guess i didn't sorry it would turn out a positve 100 not neg.

jimthompson5910 (jim_thompson5910):

now take the square root of 100 to get ?

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

so is \[\Large \sqrt{x^2} = x\] true when x is a negative number?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

good, the left side is always positive, while the right side of just x is negative (when x is negative)

jimthompson5910 (jim_thompson5910):

to fix this, we replace x on the right side with |x|

jimthompson5910 (jim_thompson5910):

so \[\Large \sqrt{x^2} = |x|\] is definitely true for all real numbers